Q. Two wheels are mounted side by side and each is marked with dot on its rim. The two dots are aligned woth the wheels at rest, then one wheel is given a constant angular acceleration of pi/2 rad/s^2 and the other pi/4 rad/sec^2. Then the two dots become aligned again for the first time after1) 2s 2) 4s3) 1s4) 8s

Grade:11

2 Answers

Yashika
5 years ago
Douse
Yashika
5 years ago
Use relative concept
Relative alpha π/2-π/4 = π/4 
Relative displacement =2π
 
Use 2nd equation of relative motion
2π=1/2×π/4×t2
t=4